# Bio-inspired reconfigurable stereo vision for robotics using omnidirectional cameras

### v1

@@ Abstract
This work introduces a novel bio-inspired reconfigurable stereo vision system for robotics, leveraging omnidirectional cameras and a novel algorithm to achieve flexible visual capabilities. Inspired by the adaptive vision of various species, our visual system addresses traditional stereo vision limitations, i.e., immutable camera alignment with narrow fields of view, by introducing a reconfigurable stereo vision system to robotics. Our key innovations include the reconfigurable stereo vision strategy that allows dynamic camera alignment, a robust depth measurement system utilizing a nonrectified geometrical method combined with a deep neural network for feature matching, and a geometrical compensation technique to enhance visual accuracy. Implemented on a metamorphic robot, this vision system demonstrates its great adaptability to various scenarios by switching its configurations of 316° monocular with 79° binocular field for fast target seeking and 242° monocular with 150° binocular field for detailed close inspection.
@@ I INTRODUCTION
To address the problems mentioned above, we turn to nature for heuristics. Present robots with bio-inspired designs have exhibited profound impacts in various intrinsic structures and biomimetic motions, including quadruped running, entangled grasping, and continuous jumping [14, 15, 16]. Furthermore, the metamorphic design is inspired by the morphology variation and performance development during the growth process of typical species. Promising progress of the metamorphic design, such as various gaits of a quadruped robot realized by changing the relative position of the joint axis [17], swimming and crawling modes switching by tuning the limbs of the turtle robot [18], and different moving strategies of the transformable origami exoskeleton [19], has been made to enhance their locomotion characteristics and adapt to versatile environments and tasks. Most of the existing bio-inspired designs focus on their mechanical structures and ignore the abundant vision features. Herbivores, e.g., rabbits, possess a panoramic FOV (360°) but narrow binocular field ($\sim$30°)[20]. In contrast, carnivores and omnivores, like cats, possess a confined FOV ($\sim$186°) but a broad binocular field ($\sim$98°) (Fig. 1(a))[21]. The differences in FOV among different species are due to the alignment of the eyes. Realizing a reconfigurable stereo vision system would benefit robots by enabling them to explore different scenarios with tunable vision features.
@@
To achieve the proposed vision system, a vision workflow algorithm that matches the system’s requirements and features is necessary. Currently, common vision algorithms rely on the epipolar constraint and pin-hole camera assumption, which requires two cameras in a stereo vision system to be fixed on the same plane with the same imaging focal length and horizontally aligned [7, 8]. The current algorithmic limitation restricts the realignment of the cameras and confines the reconfigurability of the mainstream stereo vision systems. Furthermore, the tunable FOV envisaged with a reconfigurable physical platform also proposes a new challenge for the vision algorithm.
@@
Therefore, a biomimetic reconfigurable stereo vision system is proposed that can tune the alignment of the two omnidirectional cameras to create different FOV and binocular fields through a reconfigurable platform (Fig. 1(b-c)). The vision workflow algorithm modifies the nonrectified geometrical method [22], which typically utilizes the retrieved optical path for depth measurement of the system to maintain the wide FOV. Additionally, a deep neural network [23] is employed to match local features reliably, and a geometrical compensation method is utilized to optimize and filter the results. This endows the system with the capability to collect stereo point clouds for multiple purposes, offering flexibility for a reconfigurable wide FOV vision system. Finally, the reconfigurable stereo vision system is deployed on a metamorphic robot, Origaker [24], providing this robot with various tunable vision features, including a large FOV for target-seeking and a large binocular field for detailed inspection, demonstrating the excellent adaptability for versatile applicable scenarios.
@@
2. A depth measurement system for binocular vision was developed based on the improved nonrectified stereo method with a deep neural network feature matcher and a geometrical compensation method, enabling precise depth measurement for a non-planar camera alignment.
@@ II Vision algorithm of the reconfigurable vision
To obtain a biomimetic and wide FOV, two omnidirectional cameras are employed in our system. The conventional graphic process relies on methods like rectification and epipolar alignment, which typically results in a loss of both FOV and accuracy. Additionally, the technique is not applicable to reconfigurable design. Therefore, a geometric depth measurement approach [22] is utilized to maintain the edge region of omnidirectional images for depth measurement in non-planar camera settings. However, the native implementation of this method is constrained by precision limitations, with errors propagating from several aspects, such as extrinsic calibration and homologous point matching. While the point set acquired is feasible for navigation, it remains a challenge to use as a universal visual method. To address this problem, a neural network matching method is employed to obtain reliable homologous points in complex and challenging environments robustly, and a geometric compensation method is integrated to optimize and filter the calculated spatial points. This approach enables the system to acquire a higher quality set of spatial point information, thereby enhancing its potential to be used as a general visual method. All spatial information obtained from our vision system is from this non-rectification method, including the setup of the system.
@@ II-A Optical Path Retrieve Model
With the knowledge that the incident optical path passes through the origin of the camera frame (the principal point), the optical path of an image point of interest is now retrieved for the subsequent processes, including calibration and depth computation.
@@ II-B Calibration
To perform stereo vision depth measurements, both the intrinsic and extrinsic parameters of both cameras are needed. For intrinsic calibration, each of the two omnidirectional cameras is calibrated separately using the Kannala-Brandt model with OpenCV’s built-in module [25, 26]. For extrinsic calibration, the conventional approach typically involves rectifying the omnidirectional image before performing extrinsic calibration using the usual pin-hole camera model approach. However, the existing methods prove ineffective in our setup due to the non-planar camera placement and the narrow overlapped vision, which is often located at the edge regions in certain visual modes and would be cropped in the rectification. Hence, a novel method is developed for extrinsic calibration of non-planarly placed omnidirectional cameras in our reconfigurable systems.
Using the chessboard calibration board, OpenCV’s built-in grid point detector is leveraged to accurately determine sub-pixel coordinates of the chessboard corners in the captured image. Each chessboard corner corresponds to a reverse incident optical path, which allows the system to identify a bundle of rays, each ray connecting the camera center with a grid point on the board. Utilizing the known distribution of the chessboard grid points, an optimization problem is formulated to identify a plane whose intersections with the ray bundle best fit the expected distribution. By solving this optimization problem, the 3D position of the chessboard corners could be precisely determined in the camera frame in one calibration picture.
Since our proposed method requires stereo vision overlap, the chessboard is positioned within the overlapping FOV of both cameras. This allows the system to determine the 3D spatial coordinates of the chessboard grid points in each camera frame with the optimizing solution above. A least-squares fitting method[27] is then applied to fit the relative pose between the two camera frames, minimizing the reprojection error. Using this approach, the extrinsic parameters of the cameras are acquired.
@@ II-C Homologous Point Matching
In stereo vision systems, identifying homologous points across different camera views is crucial for depth estimation and 3D reconstruction. However, due to the flexible nature of our stereo vision setup, the system cannot rely on traditional epipolar constraints and feature-based methods to find these point pairs. Instead, an advanced deep neural network structure, LightGlue[23], is employed to achieve accurate and efficient feature matching. LightGlue provides a robust solution for challenging match cases such as large changes in view angle, small visual overlaps, and challenging lighting environments with precision and efficiency, which meets the needs of our system.
@@ II-D Depth Computation Through Pseudo Intersection of Retrieved Optical Paths
The spatial location of the homologous point is calculated by triangulation of the retrieved optical path and extrinsic parameters of the camera set. Ideally, these physical points lie at the intersection of the respective optical paths on the intersection plane (determined by the two camera frame origins and the ideal optical path intersection). However, in practice, due to matching errors and image degeneration throughout the hardware and workflow, it is mathematically challenging for two optical paths to intersect precisely. To address this issue, a geometric correction mechanism is introduced that projects the two optical paths onto a predicted intersection plane to find a pseudo intersection to represent the spatial location of the homologous point (Fig. 3). The predicted intersection plane is determined by the origins of the two camera frames and the midpoint of the shortest distance between the two optical paths[28]. S is the midpoint between the two optical paths, Ll and Lr. The two projected optical paths, ll and lr intersect on the predicted plane to form the pseudo intersection S’.
@@
Figure 4: Comparison of the performance of the proposed pseudo intersection correction method with the nonrectified stereo method. The red shaded area at the center represents the slightly mismatched zone, while the yellow shaded area is for the vertical mismatch tolerance zone, and the blue shaded area is for the severely mismatched zone.
This proposed pseudo intersection correction method, showing its great monotonicity against different matching errors, has presented its advantages in both error minimizing and error filtering. To minimize errors, finding the proposed pseudo intersection provides a more precise and robust guess for noisy or mismatched inputs.
Currently, in the workflow, the main error source results from the mismatch of the homologous points, which is a common and inevitable event in the vision pipeline. The mismatched point pairs are divided into 3 zones, the slightly mismatched zones are defined for pairs where the resulting overall x-y plane error on the normalized directional vector of the corresponding optical path falls within the range of $[-0.005,0.005]$, the vertical mismatch tolerance zone is for inputs in the range $\left\{x\in[-0.005,0.005],y\in[-0.015,0.015]\right\}$ and the severely mismatched zone is the ones with error out of the two range introduced above. The inherent objective is to preserve homologous point pair inputs in the slightly mismatched and vertical mismatch tolerance zone to minimize error on computed depth while simultaneously discarding those in the severely mismatched region that do not accurately reflect the location of any spatial point.
While point pairs out of the slightly mismatched region usually mean they are illegal for depth computation, the vertical mismatch tolerance zone could still reflect reliable spatial information. Mismatches in this region have higher deviations in the optical path from the ideal intersection plane than the slightly mismatched cases. This part remains a valuable visual resource for stereo vision applications, as real-world objects, commonly upright and grounded, exhibit vertical lines in images. Due to the lack of feature points on the lines, vertical mismatches happen frequently in practice. Directly filtering out this part of homologous points would cause a non-neglectable loss in the information acquired. With our compensation method, inputs that fall in this region still hold high confidence in the computed depth, increasing the confident visual information amount. While the non-rectification method in ORB-SLAM3[22] directly uses the midpoint of the shortest distance as the intersection, making it weak against noisy input involved with this kind of mismatch
A test was performed for the nonrectified stereo method from ORB-SLAM3[22] and the proposed pseudo intersection method, as shown in Fig. 4. The test was set on a known ideal intersection at (-3000, 2000, 5000) with rays originating from (-75, 0, 0) and (75, 0, 0). The x and y values represent the noise added to the normalized direction vector of the right camera ray to simulate the mismatch and other errors. The z value represents the resulting error in depth computation in mm. The error out of [-500, 500] was filtered out for the depth output (±$\pm 10\%$).
In scenarios involving minimally mismatched pairs and predominant vertical error, the proposed method delivers heightened precision through effective error compensation. On the other hand, when confronted with severely mismatched pairs, the method’s commendable monotonicity ensures that high-error results are systematically excluded, aligning with the intended filtering criteria. In contrast, the unrectified stereo method diminishes the available visual information by prematurely filtering out inputs within the vertical mismatch tolerance zone. Furthermore, its nonuniform error distribution complicates the filtering process, as it encroaches upon the unreliable domain of severe mismatches, thereby extending into regions deemed unsuitable for accurate analysis.
@@ III-A Hardware of the Reconfigurable Vision
The testing platform of the reconfigurable vision consists of a dovetail groove gear rack and pinion translation stage (HLWX60-L250 by HENG YANG GUANG XUE) to provide the relative translation and a three-dimensional gear gimbal to provide the relative rotation of the two omnidirectional cameras (Fig. 5). The cameras’ lenses, SR1096A by SIR TEC, offer a wide image angle with horizontal (H: 195.95°±4°), vertical (V: 159.8°±3°), and diagonal (D: 199.2°±4°) FOV. The optical sensor of the camera is the CV4002 CMOS Image Sensor by CVSENS. To fully utilize the large horizontal FOV of our reconfigurable vision system, the cameras were aligned horizontally during the tests.
Figure 5: Setup of the robot vision test platform with omnidirectional cameras and tunable stages.
The system could be deployed in any relative pose as long as there is a common vision. For biomimetic purposes, three modes of the visual system and their specialties are demonstrated with our robotic vision test platform:
1) Herbivorous Vision: This mode features an exceptionally large FOV (316°) achieved by combining the view ranges of two omnidirectional cameras. It includes a relatively small angle of stereo vision for minimal 3D sensing (76°).
2) Carnivorous Vision: This mode offers wide-range stereo vision (136°) with a highly overlapped view, enhancing depth perception and precision. The total FOV of this mode is 256°.
3) Human Vision: This mode includes comprehensive depth vision characterized by a broad depth of field, with all its $\sim$196° vision field binocular.
These modes illustrate the adaptability and specialized functionalities of robotic vision systems inspired by biological features. To deploy the reconfigurable vision system, both omnidirectional cameras were calibrated for intrinsic parameters first. Then, the system underwent the working process, which included stereo extrinsic calibration, homologous points matching, and final depth computation. The results demonstrated consistency with our designed purpose and biomimetic avatars (Fig. 6(a-c)).
Figure 6: Experiment results of the 3 visual modes with depth information of each feature point dyed according to the color bar. Points with a depth value greater than 5000 mm are filtered out. (a) Herbivorous vision mode. (b) Carnivorous vision mode. (c) Human vision mode.
@@ III-B Characterization of FOV
For the constant hardware, the reconfigurability of the system is demonstrated by changing the relative spatial position of the two omnidirectional cameras. With the ultra-wide FOV of these cameras, translating them primarily adjusts the baseline length of the stereo system, thereby affecting the accuracy of homologous point matching and depth computation. In contrast, rotating the cameras plays a more significant role in determining the characteristics of different vision modes, which is achieved by balancing the size of the stereo measurement zone with the extent of the blind zone, allowing for optimal performance across various biomimetic vision applications.
@@
1) Stereoscopy Measuring Zone: The region within the overlapped FOV used for accurate depth measurement.
@@
3) Blind Zone: Areas outside the FOV of both cameras, where no visual information is available.
This spatial division is crucial for effectively utilizing the stereo vision system and is determined by the system configurations. Based on the cameras’ FOV, the relative relation of the visual fields for the proposed three zones can be determined accordingly (Fig. 7). As the stereo measurement zone (S) increases, the monocular perception zone (M) decreases at twice the rate; the blind zone (B) increases proportionally with the stereo measurement zone (S), with an offset of 32°.
Figure 7: The relation among three vision zones, stereo measurement zone (blue line), monocular perception zone (orange line), and blind zone (green line).
[…]
where S, M, and B represent the field angle of 3 zones, respectively, H represents the FOV of the omnidirectional camera used, and O represents the overlapped field angle of the binocular system.
Figure 8: The demonstration of the proposed reconfigurable vision system on the metamorphic robot Origaker. Measured depth value colored according to the color bar.In the figure, the head link configuration of Origaker and the vision system is shown, along with the actual scenarios and corresponding vision information acquired from (a) The Gecko form. (b) The Spider form. (c) The Stick-bug form.

### v2

@@ Abstract
This work introduces a novel bio-inspired reconfigurable stereo vision system for robotics, leveraging omnidirectional cameras and a novel algorithm to achieve flexible visual capabilities. Inspired by the adaptive vision of various species, our visual system addresses traditional stereo vision limitations, i.e., immutable camera alignment with narrow fields of view, by introducing a reconfigurable stereo vision system to robotics. Our key innovations include the reconfigurable stereo vision strategy that allows dynamic camera alignment, a robust depth measurement system utilizing a nonrectified geometrical method combined with a deep neural network for feature matching, and a geometrical compensation technique to enhance visual accuracy, even in very limited visual overlap ($\sim$56°). Implemented on a structural reconfigurable robot, this vision system demonstrates its great adaptability to various scenarios by switching its configurations of $\sim$207° monocular with $\sim$76° binocular field for fast target seeking and $\sim$95° monocular with $\sim$140° binocular field for detailed close inspection.
@@ I INTRODUCTION
To address the problems mentioned above, we turn to nature for heuristics. Present robots with bio-inspired designs have exhibited profound impacts in various intrinsic structures and biomimetic motions, including quadruped running, entangled grasping, and continuous jumping [14, 15, 16]. Furthermore, the metamorphic design is inspired by the morphology variation and performance development during the growth process of typical species. Promising progress of the metamorphic design, such as various gaits of a quadruped robot realized by changing the relative position of the joint axis [17], swimming and crawling modes switching by tuning the limbs of the turtle robot [18], and different moving strategies of the transformable origami exoskeleton [19], has been made to enhance their locomotion characteristics and adapt to versatile environments and tasks. Most of the existing bio-inspired designs focus on their mechanical structures and ignore the abundant vision features. Fig. 1(a) illustrated that Herbivores, e.g., rabbits, possess a panoramic FOV (360°) but narrow binocular field ($\sim$ 30°)[20]. In contrast, carnivores and omnivores, like cats, possess a confined FOV ($\sim$ 186°) but a broad binocular field ($\sim$ 98°) [21]. The differences in FOV among different species are due to the alignment of the eyes. A reconfigurable stereo vision system would benefit robots by enabling them to explore different scenarios with tunable vision features.
@@
To achieve the proposed vision system, a vision workflow algorithm that matches the system’s requirements and features is necessary. Currently, common vision algorithms rely on the epipolar constraint and the pinhole camera assumption, while also presuming that the two cameras in a stereo vision system are fixed on the same plane, have closely matched focal lengths, and are horizontally aligned.[7, 8]. The current algorithmic limitation restricts the realignment of the cameras and confines the reconfigurability of the mainstream stereo vision systems. Furthermore, the tunable FOV envisaged with a reconfigurable physical platform also proposes a new challenge for the vision algorithm.
@@
Therefore, a biomimetic reconfigurable stereo vision system is proposed that can tune the alignment of the two omnidirectional cameras to create different FOV and binocular fields through a reconfigurable platform (Fig. 1(b-c)). The vision workflow algorithm modifies the nonrectified geometrical method [22], which typically utilizes the retrieved optical path for depth measurement of the system to maintain the wide FOV. Additionally, a pre-trained deep neural network[23] is employed to match local features reliably, and a geometrical compensation method is utilized to optimize and filter the results. This endows the system with the capability to collect stereo point clouds for multiple purposes, offering flexibility for a reconfigurable wide FOV vision system. Finally, the reconfigurable stereo vision system is deployed on a metamorphic robot, Origaker [24], providing this robot with various tunable vision features, including a large FOV for target-seeking and a large binocular field for detailed inspection, demonstrating the excellent adaptability for versatile applicable scenarios.
Figure 2: The vision pipeline of the proposed system. The pipeline performs mainly three tasks: Homologous point matching, optical path retrieval, and the computation of the pseudo intersection.
@@
2. A depth measurement system for binocular vision was developed based on the improved nonrectified stereo method with a pre-trained deep neural network feature matcher and a geometrical compensation method, enabling precise depth measurement for a non-planar camera alignment.
@@ II Vision algorithm of the reconfigurable vision
To obtain a biomimetic and wide FOV, two omnidirectional cameras are employed in our system. The conventional graphic process relies on methods like rectification, which typically results in a loss of both FOV and accuracy. Additionally, the technique is not applicable to reconfigurable design. Therefore, a geometric depth measurement approach [22] is utilized to maintain the edge region of omnidirectional images for depth measurement in non-planar camera settings. However, the native implementation of this method is constrained by precision limitations, with errors propagating from several aspects, such as extrinsic calibration and homologous point matching. While the point set acquired is feasible for navigation, it remains a challenge to use as a universal visual method. To address this problem, an off-the-shelf neural network module is employed to obtain reliable homologous points in complex and challenging environments robustly, and a geometric compensation method is integrated to optimize and filter the calculated spatial points. This approach enables the system to acquire a higher quality set of spatial point information, thereby enhancing its potential to be used as a general visual method. All spatial information obtained from our vision system is from this nonrectification method, including the setup of the system. Fig. 2 depicts the functional component of the vision pipeline.
@@ II-A Optical Path Retrieve Model
With the knowledge that the incident optical path passes through the origin of the camera frame, the optical path of an image point of interest is now retrieved for the subsequent processes, including calibration and depth computation.
@@ II-B Calibration
To perform stereo vision depth measurements, both the intrinsic and extrinsic parameters of both cameras are needed. For intrinsic calibration, each of the two omnidirectional cameras is calibrated separately using the Kannala-Brandt model with OpenCV’s built-in library [25, 26]. For extrinsic calibration, the conventional approach typically involves rectifying the omnidirectional image before performing extrinsic calibration using the usual pin-hole camera model approach. However, the existing methods prove ineffective in our setup due to the non-planar camera placement and the narrow overlapped vision, which is often located at the edge regions in certain visual modes and would be cropped in the rectification. Hence, a novel method is developed for the extrinsic calibration of non-planarly placed omnidirectional cameras with possibly large camera viewpoint disparity in our reconfigurable systems.
Using the chessboard calibration board, OpenCV’s built-in grid point detector is leveraged to accurately determine the sub-pixel coordinates of the chessboard corners in the captured image. Each chessboard corner corresponds to a reverse incident optical path, which allows the system to identify a bundle of rays, each ray connecting the camera center with a grid point on the board. Utilizing the known distribution of the chessboard grid points, an optimization problem is formulated to identify a plane whose intersections with the ray bundle best fit the expected distribution. By solving this optimization problem, the 3D position of the chessboard corners could be precisely determined in the camera frame in one calibration picture.
Since our proposed method requires stereo vision overlap, the chessboard is positioned within the overlapping FOV of both cameras. This allows the system to determine the 3D spatial coordinates of the checkerboard grid points in each camera frame with the optimizing solution above. A least-squares fitting method[27] is then applied to fit the relative pose between the two camera frames, minimizing the reprojection error. Using this approach, the extrinsic parameters of the cameras are acquired.
@@ II-C Homologous Point Matching
In stereo vision systems, the identification of homologous points across different camera views is a critical step for accurate depth estimation and 3D reconstruction. However, despite recent advancements in stereo-matching techniques, conventional methods have proven inadequate for our flexible system, which requires robust homologous point matching in unrectified fisheye images with potentially extreme viewpoint variations. Traditional feature-based methods, such as SGBM and SIFT, are ill-suited for this task due to their inability to handle significant lens distortion and large angular separations between cameras, which are inherent in our setup. Similarly, state-of-the-art neural network-based stereo approaches, including DISPNET and RAFT-STEREO, have also fallen short. These methods are predominantly trained on datasets captured with pinhole cameras and small viewpoint disparities, and many of them output disparity maps rather than direct homologous point pairs, making them incompatible with our optical path methodology. To address these challenges, we employ [23], an advanced deep neural network architecture for local feature matching, leveraging its pre-trained weights to provide a robust and efficient solution for demanding conditions such as large viewpoint changes, limited visual overlap, and challenging lighting environments. Without additional training or fine-tuning, LightGlue enables accurate and reliable homologous point identification in complex scenarios, overcoming the limitations of traditional and neural network-based stereo-matching methods. This is particularly significant for applications involving fisheye cameras and large baseline configurations, where conventional techniques often fail to deliver satisfactory results. A comparison of common used SGBM algorithm and LightGlue’s matching result is displayed in Fig. 4, where SGBM algorithm produces sparse matches with a notable fraction of errors. In contrast, LightGlue generates a denser set of matching pairs with significantly improved accuracy.
@@ II-D Depth Computation Through Pseudo Intersection of Retrieved Optical Paths
The spatial location of the homologous point is calculated by triangulation of the retrieved optical path and extrinsic parameters of the camera set. Ideally, these physical points lie at the intersection of the respective optical paths on the intersection plane (determined by the two camera frame origins and the ideal optical path intersection). However, in practice, due to matching errors and image degeneration throughout the hardware and workflow, it is mathematically challenging for two optical paths to intersect precisely. To address this issue, a geometric correction mechanism is introduced that projects the two optical paths onto a predicted intersection plane to find a pseudo intersection to represent the spatial location of the homologous point (Fig. 5). The predicted intersection plane is determined by the origins of the two camera frames and the midpoint of the shortest distance between the two optical paths[28]. $S$ is the midpoint between the two optical paths, $L_{l}$ and $L_{r}$. The two projected optical paths, $l_{l}$ and $l_{r}$ intersect on the predicted plane to form the pseudo intersection $S^{\prime}$.
Figure 4: Stereo matching results of (a) SGBM algorithm and (b) pre-trained LightGlue.
@@
This proposed pseudo intersection correction method, showing its great monotonicity against different matching errors, has presented its advantages in both error minimizing and error filtering. To minimize errors, finding the proposed pseudo intersection provides a more precise and robust guess for noisy or mismatched inputs. In the current workflow, the primary source of error arises from mismatched homologous points, which are common in the vision pipeline, especially in the presence of distortion and large camera angle separations. Vertical mismatches, often occurring due to the lack of feature points on vertical lines, are particularly frequent in practice. However, vertical mismatches remain a valuable visual resource for stereo vision applications, as real-world objects, which are typically upright and grounded, produce vertical lines in images. Unlike the original nonrectification method, which discards points with vertical mismatches, the proposed method retains this critical information, improving depth estimation accuracy.
@@ III-A Hardware of the Reconfigurable Vision
The testing platform of the reconfigurable vision consists of a linear stage and a pinion translation stage to provide the relative translation and rotation of the two omnidirectional cameras (Fig. 6). The lenses of the cameras, SR1096A from SIR TEC, offer a wide angle of view with horizontal (H: 195.95°±4°), vertical (V: 159.8°±3°), and diagonal (D: 199.2°±4°) FOV. The optical sensor of the camera is the CV4002 CMOS Image Sensor by CVSENS. To fully utilize the large horizontal FOV of our reconfigurable vision system, the cameras were horizontally aligned during the tests.
Figure 6: Hardware setup of the robot vision test platform with omnidirectional cameras and tunable stages.
To deploy the reconfigurable vision system, intrinsic parameters for both omnidirectional cameras were first calibrated. Subsequently, the system underwent a series of processes, including stereo extrinsic calibration, homologous point matching, and final depth computation. The results showed consistency with our intended design and biomimetic avatars.
@@ III-B Characterization of FOV
The reconfigurability of the system, with constant hardware, is demonstrated by varying the relative spatial positions of the two omnidirectional cameras. Due to the ultra-wide FOV of these cameras, translating them mainly adjusts the stereo system’s baseline length, which in turn affects the accuracy of homologous point matching and depth computation. In contrast, rotating the cameras has a more profound impact on the characteristics of various vision modes, balancing the stereo measurement zone with the blind zone size, thus ensuring optimal performance in different biomimetic vision applications.
@@
1) Stereo Measuring Zone: The region within the overlapped FOV used for accurate depth measurement.
@@
3) Blind Zone: Areas outside the FOV of both cameras, where visual information is not available.
This spatial division is crucial for effectively utilizing the stereo vision system and is primarily determined by the system configurations. Defining the field angle of each vision zone is also essential in evaluating the system’s perceptual capabilities. A common approach is to determine the corresponding FOV width relative to different object plane positions. However, since the field angle of the vision zones in the proposed system may exceed 180°, this conventional method is not applicable. To address this limitation, we introduce the concept of the equivalent field angle to better define the system’s FOV, as illustrated in Fig. 7 (a).
This concept abstracts the binocular camera system as a single viewpoint located at the midpoint of the stereo baseline (assuming a symmetrical configuration). Given an equivalent object distance $R$ for this viewpoint, the perception zones can be represented as arcs on a circle with a radius $R$. The equivalent field angle of each zone is defined as the central angle corresponding to the arc. The three vision zones’ equivalent field angle $\theta_{stereo}$, $\theta_{monocular}$, $\theta_{blind}$ could be derived with known camera FOV $\theta_{FOV}$, baseline length $B$, camera main axis separation angle $\theta_{separation}$ and object distance $R$:
$\left\{\begin{aligned} \theta_{stereo}&=u\left[1-\frac{B}{2R}\cot\!\left(\frac{u}{2}\right)\right]\\ \theta_{blind}&=u+2\,\arcsin\!\left[\frac{B}{2R}\,\sin\!\left(\frac{\pi+u}{2}\right)\right]\\ \theta_{monocular}&=2\pi-\theta_{stereo}-\theta_{blind}\end{aligned}\right$. (3)
Figure 7: Equivalent field angle. (a) Perception zones and corresponding equivalent field angle. (b) Variation of stereoscopy measuring the zone’s equivalent field angle $\theta_{stereo}$ along with separation angle $\theta_{separation}$ and equivalent object angle $R$. (c) Variation of stereoscopy measuring the zone’s equivalent field angle $\theta_{monocular}$ along with separation angle $\theta_{separation}$ and equivalent object angle $R$.
where
$u=\theta_{FOV}-\theta_{separation}$ (4)
With the complex possible configurations of the system, Eq. (3) expresses the equivalent field angle of three perception zones in the general circumstance ($\theta_{separation}\geq\theta_{FOV}-\pi$, $R\geq\max\!\left(\frac{B}{2}\tan\!\left(\frac{u}{2}\right),\,\frac{B}{2}\right)$.) Fig. 7 (b) and (c) serves as a visualization for the variation of $\theta_{stereo}$ and $\theta_{monocular}$ with respect to separation angle $\theta_{separation}$ and equivalent object angle $R)$. With smaller $\theta_{separation}$, the binocular camera tends to have more vision overlap, leading to a larger depth perception field, and vice versa.
III-C Characterization of Measurement Precision
With the biomimetic and reconfigurable design strategy, the system can be deployed in any relative pose, provided that a common vision field exists. To evaluate the measurement precision of our system, two typical modes of the visual system and their unique characteristics are demonstrated using our vision test platform, as displayed in Fig. 8:
Figure 8: Reconfigurable vision precision testing experiment setup. (a) Herbivorous vision mode. (b) Human vision mode.
a) Herbivorous Vision mode: This mode features an exceptionally large FOV ($\sim$265°) achieved by combining the view ranges of two omnidirectional cameras. It includes a relatively small angle of stereo vision for minimal 3D sensing ($\sim$56°). By testing on this configuration, the system’s ability for depth measurement on the lateral region of the omnidirectional camera is demonstrated.
b) Human Vision mode: This mode includes comprehensive depth vision characterized by a broad depth field, with a $\sim$193° binocular vision field. Through testing on this configuration, the system’s ability to perform wide-range depth sensing is validated.
The field angle value presented here is calculated based on the actual experimental setup, with the equivalent object distance $R$ set to 350 mm. A checkerboard target (20 mm $\times$ 20 mm $\times$ 350 mm), along with its holder, is mounted on an optical breadboard alongside the vision system, and its grid points are measured using different vision approaches. In human vision mode, where the conventional rectified extrinsic calibration method remains applicable due to the wide vision overlap, three measurement processes are conducted:
1) The conventional rectified binocular method: The camera’s extrinsic parameters are estimated through rectified calibration pictures, and the measured grid points are also found on rectified pictures. The spatial measurement assumes the epipolar constraint is satisfied during the rectification process and directly finds the disparity through the difference of $x$ coordinates of the grid points. The depth $Z$ is calculated by:
$Z=\frac{f\cdot B}{d}$ (5)
where $f$ represents the focal length, $B$ represents the length of baseline, and $d$ represents the pixel disparity. The 3D location is found by reprojecting the 2D points to 3D space.
2) The nonrectified optical path method: With extrinsic parameters still estimated through a rectified calibration picture set, triangulation is performed using the retrieved optical path. The intersection is simply computed as the midpoint of the closest distance between the two incident optical paths.
3) The proposed geometrical-compensation nonrectified optical path method: The proposed method uses the nonrectified calibration method introduced in Section II.BII.B for extrinsic parameters and uses the pseudo intersection method introduced in Section II.DII.D for the 3D measurement.
For the herbivorous vision mode, which contains a limited vision overlap ($\sim$56°), the conventional extrinsic calibration method fails because, after rectification, the checkerboard gets cropped in the process, preventing successful corner detection. Thus, only the measurement result of the proposed method is shown, as the other two methods cannot be performed without extrinsic parameters.
The depth measurement results of the two modes are shown in Fig. 9(a)–(d) in the following order: human mode conventional binocular method, human mode non-rectified method, human mode proposed method, and herbivorous mode proposed method. The absolute errors of depth measurement are 0.793 mm, 2.36 mm, 1.08 mm, and 1.44 mm, respectively. The relative errors are 0.227%, 0.675%, 0.308%, and 0.413%, respectively. The experimental results demonstrate the system’s high measurement precision while preserving a wide FOV without rectification, even in extreme conditions such as narrow visual overlaps ($\sim$56°). The equivalent field angle of perception zones in all four test groups is shown in Fig. 9(e).
Figure 9: Depth Measurement value and error of grid point with different methods. (a) Measurement. (b) Absolute error. (c) Relative error. (d) Standard deviation. (e) Equivalent field angle profile of the three perception fields in all corresponding methods.
Therefore, our proposed method achieves measurement precision comparable to current baseline methods, even in extreme conditions where depth must be detected within the lateral regions of the omnidirectional image. The high precision further validates the system’s capacity to fully exploit its entire camera FOV for depth sensing, supporting the biomimetic principle that all overlapping vision fields can be used for stereo measurement, thereby justifying the proposed system’s extremely wide FOV. Meanwhile, with acceptable measurement precision, the generally wider FOV also demonstrated the improved measurement ability.